OpenStudy (anonymous):

a population of 1750 cheetahs decreases by 11% per year. how many cheetahs will there be in the population after 10 years? round to the nearest whole number.

OpenStudy (anonymous):

1. find 11% of the total number of the cheetahs which is 1750 2. multiply by 10 3. subtract from the original

OpenStudy (anonymous):

ok I am still not understanding

OpenStudy (anonymous):

She's telling you to: 1750 * 11% = X. Take X * 10 = X. Take that X and subtract it from 1750

OpenStudy (zehanz):

Another way to do it: look at what is left after one year: 100%-11%=89%. This means, because 89% actually means: 89/100 = .89, you can multiply 1750 with 0.89 to get the number of cheetahs after one year. Next year: again multiply with 0.89, or (0.89)² in total. After 10 years: 1750*(0.89)^10

OpenStudy (zehanz):

You have to take into account that each year the population decreases with 11%. The first year, that is 11% of 1750. The next year, it is 11% of the population in that year (which is smaller than 1750!). It is exponential "growth"...